OpenStudy (anonymous):

Hi, I need help with this Algebra question please. It's Subtract 3x² - 2xy + 7y² from -8x²-5xy+9y²

OpenStudy (anonymous):

Things with simmilar variables may be subtracted from eachother so, 3x^2 - (-8x^2) would be 11x^2 (remember, minus a - is a positive) keep on doing that for all the rest, and you should get the correct answer.

OpenStudy (lgbasallote):

@nkbxwb hmmm seems you got it mixed up... 3x^2 is subtracted FROM -8x^2 so it's -8x^2 - 3x^2 and so on.. ^_^

OpenStudy (anonymous):

if you do this websit for that and ever thing you well do fine mathwayslovetheequation.com

OpenStudy (anonymous):

@lgbasallote oops! It's late and its a Monday! aha

OpenStudy (anonymous):

Please help me with this question. Subtract the sum of 2a ²- 3ab + 4b ² and 6a² + 2ab -2b² from 3a²- 2b² And can you please provide to me an explanation of the steps of what you did first, second and so on? It would help so much.

OpenStudy (anonymous):

Ok, for your second question You must combine the like terms of the first 2 equations, 2a^2-3ab+4b^2 +6a^2+2ab-2b^2 Answer for that is, 8a^2-ab+2b^2 Then you have to subtract terms of the previous equation with 3a^2-2b^2 8a^2-ab+2b^2 -3a^2 -2b^2 <--- no -ab in the bottom equation just carry it down Answer is, 5a^2-ab (You should check if I did my math correctly!)

OpenStudy (anonymous):

Second opinion please?

OpenStudy (anonymous):

http://www.freemathhelp.com/ it will give you the answer

OpenStudy (lgbasallote):

my second opinion would be that it should be 3a^2 -2b^2 -(8a^2 - ab + 2b^2) because the sum is subtracted FROM...

OpenStudy (anonymous):

I need help with this question, please, X ²y (3x²- 2xy + 2y²) Thank you, Can you please write the steps of how you arrived to the conclusion? I sincerely appreciate it.

OpenStudy (lgbasallote):

x^2 y? is that the first term?

OpenStudy (anonymous):

So I've been trying to use the freemathhelp.com website And apparently it's still not working. I inserted this equation... X ²y (3x²- 2xy + 2y²) But when I hit enter, the message appearing on my screen is that it doesn't recognize the "square" I tried it with this symbol "^" that didn't work there, what am I doing wrong?

OpenStudy (anonymous):

I have no idea what you're doing wrong because when I enter it I get an answer for it. I don't feel like copying it over is going to help you learn anything

OpenStudy (anonymous):

are you in the algebra tab? cause if not that could be why it's not working

OpenStudy (anonymous):

I need help with 2 radical 2over 2 radical 5 subtracted by radical 2\[2\sqrt{2} \over 2\sqrt{5} -\sqrt{2}\]

OpenStudy (anonymous):

The directions are express in simplest radical form.

OpenStudy (anonymous):

The second problem I have is